OpenStudy (anonymous):

If 8.74 g of CuNO3 is dissolved in water to make a 0.700 M solution, what is the volume of the solution ???

OpenStudy (australopithecus):

Molarity = Moles/Liters Basic Algebra Liters = Moles/Molarity

OpenStudy (abmon98):

Number of moles=Mass(g)/Molar Mass(g/mol)=8.74/125.55 Number of moles=Concentration(mol/dm^3)*Volume(dm^3)-->Volume=Number of moles/Concentration(mol/dm^3)=8.74/125.55/0.700